# MLLM-based Discovery of Intrinsic Coordinates and Governing Equations from High-Dimensional Data

### v1

@@ 4.3 System with Pixel Coordinate
4.4 System with Latent Coordinate
The reasoning results of VER in the latent coordinate system are presented in Figure 3, including the zero-shot inferred latent coordinate space and governing equations. VER takes observation sequences in the form of multi-channel image modalities as input and locates the intrinsic dimension of high-dimensional nonlinear partial differential equations. The results in Figure 3 show that, despite the high nonlinearity of the original equations, VER reveals the linear approximation of oscillatory patterns in the latent space spanned by the intrinsic dimensions, which is consistent with the conclusions of previous works [29]. Notably, in the case of the Water system, observed sequences evolved from different initial conditions are identified as trajectories with consistent dynamics in the latent space. While VER determines that two latent coordinates are required to capture the oscillatory patterns in the observation data, the results of the equation inference module indicate that the system’s dynamics are actually governed by a one-dimensional linear ODE. The second dimension is only used to distinguish coordinate region differences caused by varying initial conditions.
@@ 4.4 System with Latent Coordinate
4.5 Robustness
We add Gaussian noise with varying strengths $\sigma$ (standard deviation, mean of 00) to the observations, which biases derivative estimation and affects the accuracy and simplicity of the discovered equations. Table 6 shows the performance of all methods on the latent coordinate system. VER exhibits significantly greater robustness to noise compared to the baseline. This is because MLLM, guided by prompts, balances accuracy and equation simplicity during identification, avoiding coefficient estimation bias (overfitting) for irrelevant symbolic terms caused by noise disturbances.
4.6 Ablation Study
@@
We conduct ablation studies to validate the contributions of VER’s replaceable components to MLLM. First, we assess MLLM’s ability to infer equations directly from raw motion videos. Then, we test various ablation versions without specific modules or tools. The ablation models are evaluated on linear and cubic systems, as shown in Table 2. Results indicate that without the reasoning framework, MLLM (GPT-4o) struggles to accurately discover equations from videos, leading to significant discrepancies in reconstructed versus predicted trajectories. This shows VER is designed to unlock MLLM’s visual understanding and reasoning capabilities for equation discovery, not relying on potential data leakage during MLLM training. Additionally, equations discovered by VER-SINDy contain more redundant terms compared to VER, resulting in larger cumulative errors in long-term predictions. Thus, VER’s architecture is crucial for enabling MLLM to discover symbolic equations from high-dimensional data. Finally, the ablation version with enhanced visual prompts shows significant performance degradation compared to the full VER, confirming their importance. In addition, we conduct further ablation experiments on the alternative parameter selection modules, filter parameters (appendix B.3) and SINDy symbolic libraries (Figure 5), within the VER framework, and the results validate the superiority of MLLM compared to traditional Bayesian methods.
@@
Table 2: Ablation studies on linear and cubic systems: ’w/o’ means ’without’. VER-SINDy uses the standard SINDy algorithm instead of VER’s inference module. $R^{2}$ scale is $10^{-1}$.
Case Method Equation Prediction
Terms Found False Positives $R^{2}$ $R^{2}@100$ $R^{2}@1000$
Linear GPT-4o Yes $2.40\pm 1.36$ $2.46\pm 2.31$ $2.99\pm 2.73$ $0.67\pm 3.83$
w/o measure No $1.02\pm 0.54$ $8.18\pm 0.60$ $7.99\pm 0.53$ $6.97\pm 0.86$
w/o amplifier Yes $0.64\pm 0.31$ $9.07\pm 0.27$ $8.99\pm 0.33$ $8.82\pm 0.37$
w/o replayer Yes $0.51\pm 0.18$ $9.19\pm 0.09$ $9.26\pm 0.19$ $9.12\pm 0.21$
VER-SINDy Yes $1.25\pm 0.65$ $\mathbf{9.69\pm 0.23}$ $8.54\pm 0.69$ $8.27\pm 0.74$
VER Yes $\mathbf{0.38\pm 0.22}$ $9.59\pm 0.25$ $\mathbf{9.68\pm 0.20}$ $\mathbf{9.62\pm 0.23}$
Cubic GPT-4o No $4.40\pm 1.39$ $1.16\pm 2.71$ $1.09\pm 2.46$ $0.63\pm 2.77$
w/o measure No $2.49\pm 0.82$ $8.73\pm 0.89$ $7.35\pm 1.50$ $2.96\pm 2.20$
w/o amplifier No $1.78\pm 0.62$ $9.10\pm 0.53$ $7.63\pm 1.72$ $4.87\pm 2.59$
w/o replayer Yes $1.67\pm 0.56$ $9.26\pm 0.41$ $7.45\pm 1.43$ $5.57\pm 2.32$
VER-SINDy No $2.85\pm 0.93$ $8.79\pm 0.65$ $7.10\pm 1.71$ $3.18\pm 2.62$
5 Related Work
5.1 Inferring physics from video
@@
Many studies focus on modeling video dynamics for predictive understanding, often using parameterized modules for implicit physics. Vincent et al. [19] decoupled video content into PDE dynamics and unknowns, predicting dynamics via a parameterized PDE network. Wu et al. [20] applied stochastic PDE dynamics for real-world randomness. Hofherr et al. [22] and Bounou et al. [23] used neural ODEs and Koopman operators as dynamic predictors for interpretable parameters. These methods limit interpretability by parameterizing dynamics. With known equations, Jaques et al. [28] used inverse graphics to capture physical quantities, while Kandukuri et al. [27] enforced rigid body mechanics via convolutional encoders to identify properties like mass and friction. Yang et al. [25] and Fotiadis et al. [26] utilized encoder-decoder models for low-dimensional physical coordinates. Unlike our approach, these methods depend on prior knowledge or known equations, lacking the ability to infer physical coordinates and governing equations from unfamiliar videos.
5.2 Multimodal large language models on vision tasks
@@
Recent research on multimodal large language models in vision tasks has demonstrated significant advancements. Wu et al. [34] meticulously designed a reasoning chain to unlock the potential of MLLMs in object detection tasks. Liu et al. [12] further combined physics engines with MLLMs, enabling MLLMs to analyze motion targets and their physical properties from images to generate physics-grounded videos. Additionally, extensive research has focused on fine-tuning or reasoning chains to enhance MLLMs’ video understanding capabilities [55, 56, 57]. In contrast to these efforts, we pioneer the exploration of unlocking MLLMs’ ability to discover equations from videos through carefully designed tools and reasoning chains.
5.3 Visual symbolic regression
@@
High-dimensional symbolic regression methods for videos is a novel scientific task that typically follows two paradigms: (1) object tracking and equation discovery, where trajectories are extracted and equations identified [14, 30]; (2) encoder-decoder models for latent physical coordinates, mapping data into low-dimensional spaces for equation regression [53, 29]. Our approach uses MLLMs’ visual understanding and scientific knowledge for zero-shot discovery of both coordinates and equations, unifying these paradigms in one framework.
6 Conclusion
@@
Limitation & future work. The visual perception depends on high-quality prompts, and smoothing filters can distort high-frequency trajectories. Future efforts will refine the prompt toolkit and expand to complex systems like materials science and biology.
@@ Appendix A Data Generation
    $\partial_{t}h+\nabla\cdot h\mathbf{u}$ $=0,$ (5)
@@
    We use open-source simulation data from the related work [51].
@@ Appendix B Additional Results
B.1 Real-world motion video
@@
Resolution $500\times 500$ $500\times 500$ $500\times 500$ $500\times 500$ $500\times 500$ $500\times 500$ $100\times 100$ $64\times 64$ $128\times 128$ $128\times 128$ $128\times 128$
@@
B.2 Adaptive regularization coefficient
@@
Case Method Equation Prediction
Terms Found False Positives $R^{2}$ $R^{2}@100$ $R^{2}@1000$
Linear VER Yes $0.38\pm 0.22$ $\mathbf{9.59\pm 0.25}$ $9.68\pm 0.20$ $9.62\pm 0.23$
$\eta$-VER Yes $\mathbf{0.35\pm 0.19}$ $9.54\pm 0.20$ $\mathbf{9.70\pm 0.18}$ $\mathbf{9.64\pm 0.22}$
Cubic VER Yes $1.41\pm 0.73$ $9.53\pm 0.32$ $7.96\pm 1.60$ $6.16\pm 2.78$
Circular VER Yes 0 $9.91\pm 0.07$ $\mathbf{9.99\pm 0.00}$ $9.67\pm 0.45$
$\eta$-VER Yes 0 $\mathbf{9.92\pm 0.06}$ $9.97\pm 0.04$ $\mathbf{9.70\pm 0.42}$
VDP VER Yes $0.80\pm 0.40$ $\mathbf{9.69\pm 0.08}$ $9.35\pm 0.19$ $7.28\pm 0.72$
$\eta$-VER Yes $\mathbf{0.51\pm 0.27}$ $9.45\pm 0.05$ $\mathbf{9.52\pm 0.11}$ $\mathbf{8.31\pm 0.60}$
B.3 Bayesian hyperparameter search
@@
Case Method Equation Prediction
Terms Found False Positives $R^{2}$ $R^{2}@100$ $R^{2}@1000$
Linear Bayes-SINDy Yes $0.49\pm 0.18$ $9.37\pm 0.41$ $9.42\pm 0.35$ $9.08\pm 0.47$
Bayes-smooth Yes $2.51\pm 1.02$ $8.71\pm 0.44$ $8.56\pm 0.54$ $4.80\pm 1.22$
Cubic Bayes-SINDy Yes $3.64\pm 1.39$ $9.09\pm 0.20$ $7.51\pm 1.48$ $3.63\pm 2.11$
Bayes-smooth Yes $4.36\pm 1.96$ $7.72\pm 1.04$ $7.02\pm 1.69$ $2.94\pm 1.86$
Glider Bayes-SINDy Yes $1.05\pm 0.35$ $\mathbf{9.90\pm 0.23}$ $9.51\pm 0.14$ $8.33\pm 0.66$
Bayes-smooth Yes $1.15\pm 0.30$ $9.15\pm 0.14$ $9.29\pm 0.09$ $7.89\pm 0.26$
VER Yes $\mathbf{0.84\pm 0.17}$ $9.78\pm 0.08$ $\mathbf{9.90\pm 0.07}$ $\mathbf{9.23\pm 0.18}$
B.4 Time cost
@@ Appendix C Supplementary Materials
Bruss AE-SINDy $\mathbf{8.62\pm 0.23}$ $7.72\pm 0.44$ $6.91\pm 0.81$ $6.63\pm 0.92$
VER $8.50\pm 0.07$ $\mathbf{8.32\pm 0.18}$ $\mathbf{7.68\pm 0.61}$ $\mathbf{6.96\pm 0.82}$
Water AE-SINDy $9.64\pm 0.08$ $9.00\pm 0.35$ $8.07\pm 0.55$ $7.03\pm 1.30$

### v2

@@ Introduction
System with Latent Coordinate
The reasoning results of VER in the latent coordinate system are presented in Figure 3, including the zero-shot inferred latent coordinate space and governing equations. VER takes observation sequences in the form of multi-channel image modalities as input and locates the intrinsic dimension of high-dimensional nonlinear partial differential equations. The results in Figure 3 show that, despite the high nonlinearity of the original equations, VER reveals the linear approximation of oscillatory patterns in the latent space spanned by the intrinsic dimensions, which is consistent with the conclusions of previous works (Champion et al. 2019). Notably, in the case of the Water system, observed sequences evolved from different initial conditions are identified as trajectories with consistent dynamics in the latent space. While VER determines that two latent coordinates are required to capture the oscillatory patterns in the observation data, the results of the equation inference module indicate that the system’s dynamics are actually governed by a one-dimensional linear ODE. The second dimension is only used to distinguish coordinate region differences caused by varying initial conditions.
@@
Robustness
We add Gaussian noise with varying strengths $\sigma$ (standard deviation, mean of 0) to the observations, which biases derivative estimation and affects the accuracy and simplicity of the discovered equations. Table 6 shows the performance of all methods on the latent coordinate system. VER exhibits significantly greater robustness to noise compared to the baseline. This is because MLLM, guided by prompts, balances accuracy and equation simplicity during identification, avoiding coefficient estimation bias (overfitting) for irrelevant symbolic terms caused by noise disturbances.
Ablation Study
We conduct ablation studies to validate the contributions of VER’s replaceable components to MLLM. As shown in Table 2, we evaluate several ablated versions on linear and cubic systems. Without our reasoning framework, the base MLLM (GPT-4o) fails to discover accurate equations from raw video, leading to large trajectory prediction errors. This confirms VER unlocks the MLLM’s intrinsic reasoning for this task, rather than relying on potential data leakage. Ablating key components like the enhance visual prompts or the symbolic refiner (VER-SINDy) also lead to significant performance degradation or more redundant equation terms. Thus, VER’s architecture is crucial for discovering symbolic equations from high-dimensional data. Further ablations on parameter selection (Appendix B) and symbolic libraries (Fig. 5) confirm the MLLM’s superiority over traditional Bayesian methods. Finally, to test dependency on the backbone MLLM, we swap to Gemini-1.5-flash (Team et al. 2024). Although this weaker model reduces performance, it still successfully identifies the correct equation terms.
@@
Method Equation Prediction
Terms
Found
False
Positives
$R^{2}$ $R^{2}@100$ $R^{2}@1000$
Linear
GPT-4o Yes 2.40±1.362.40$\pm$ 1.362.40 ± 1.36 0.25±0.230.25$\pm$ 0.230.25 ± 0.23 0.30±0.270.30$\pm$ 0.270.30 ± 0.27 0.07±0.380.07$\pm$ 0.380.07 ± 0.38
w/o measure No 1.02±0.541.02$\pm$ 0.541.02 ± 0.54 0.82±0.060.82$\pm$ 0.060.82 ± 0.06 0.81±0.050.81$\pm$ 0.050.81 ± 0.05 0.70±0.090.70$\pm$ 0.090.70 ± 0.09
w/o amplifier Yes 0.64±0.310.64$\pm$ 0.310.64 ± 0.31 0.91±0.030.91$\pm$ 0.030.91 ± 0.03 0.90±0.030.90$\pm$ 0.030.90 ± 0.03 0.88±0.040.88$\pm$ 0.040.88 ± 0.04
w/o replayer Yes 0.51±0.180.51$\pm$ 0.180.51 ± 0.18 0.92±0.010.92$\pm$ 0.010.92 ± 0.01 0.93±0.020.93$\pm$ 0.020.93 ± 0.02 0.91±0.020.91$\pm$ 0.020.91 ± 0.02
VER-SINDy Yes 1.25±0.651.25$\pm$ 0.651.25 ± 0.65 $\mathbf{0.97\pm 0.02}$ 0.85±0.070.85$\pm$ 0.070.85 ± 0.07 0.83±0.070.83$\pm$ 0.070.83 ± 0.07
VER(Gemini) Yes 0.58±0.230.58$\pm$ 0.230.58 ± 0.23 0.91±0.020.91$\pm$ 0.020.91 ± 0.02 0.92±0.030.92$\pm$ 0.030.92 ± 0.03 0.89±0.050.89$\pm$ 0.050.89 ± 0.05
VER Yes $\mathbf{0.38\pm 0.22}$ 0.96±0.030.96$\pm$ 0.030.96 ± 0.03 $\mathbf{0.97\pm 0.02}$ $\mathbf{0.96\pm 0.02}$
Cubic
GPT-4o No 4.40±1.394.40$\pm$ 1.394.40 ± 1.39 0.12±0.270.12$\pm$ 0.270.12 ± 0.27 0.11±0.250.11$\pm$ 0.250.11 ± 0.25 0.06±0.280.06$\pm$ 0.280.06 ± 0.28
w/o measure No 2.49±0.822.49$\pm$ 0.822.49 ± 0.82 0.87±0.090.87$\pm$ 0.090.87 ± 0.09 0.74±0.150.74$\pm$ 0.150.74 ± 0.15 0.30±0.220.30$\pm$ 0.220.30 ± 0.22
w/o amplifier No 1.78±0.621.78$\pm$ 0.621.78 ± 0.62 0.91±0.050.91$\pm$ 0.050.91 ± 0.05 0.76±0.170.76$\pm$ 0.170.76 ± 0.17 0.49±0.260.49$\pm$ 0.260.49 ± 0.26
w/o replayer Yes 1.67±0.561.67$\pm$ 0.561.67 ± 0.56 0.93±0.040.93$\pm$ 0.040.93 ± 0.04 0.75±0.140.75$\pm$ 0.140.75 ± 0.14 0.56±0.230.56$\pm$ 0.230.56 ± 0.23
VER-SINDy No 2.85±0.932.85$\pm$ 0.932.85 ± 0.93 0.88±0.070.88$\pm$ 0.070.88 ± 0.07 0.71±0.170.71$\pm$ 0.170.71 ± 0.17 0.32±0.260.32$\pm$ 0.260.32 ± 0.26
VER(Gemini) Yes 1.96±0.741.96$\pm$ 0.741.96 ± 0.74 0.90±0.070.90$\pm$ 0.070.90 ± 0.07 0.74±0.190.74$\pm$ 0.190.74 ± 0.19 0.41±0.250.41$\pm$ 0.250.41 ± 0.25
VER Yes $\mathbf{1.41\pm 0.73}$ $\mathbf{0.95\pm 0.03}$ $\mathbf{0.80\pm 0.16}$ $\mathbf{0.62\pm 0.28}$
@@
Table 2: Ablation studies on linear and cubic systems: ’w/o’ means ’without’. VER-SINDy uses the standard SINDy algorithm instead of VER’s inference module. VER(Gemini) uses Gemini-1.5-Flash as backbone MLLM.
Related Work
Inferring physics from video
@@
Many studies focus on modeling video dynamics for predictive understanding, often using parameterized modules for implicit physics. Vincent et al. (Guen and Thome 2020) decoupled video content into PDE dynamics and unknowns, predicting dynamics via a parameterized PDE network. Wu et al. (Wu et al. 2023) applied stochastic PDE dynamics for real-world randomness. Hofherr et al. (Hofherr et al. 2023) and Bounou et al. (Bounou, Ponce, and Carpentier 2021) used neural ODEs and Koopman operators as dynamic predictors for interpretable parameters. These methods limit interpretability by parameterizing dynamics. With known equations, Jaques et al. (Jaques, Burke, and Hospedales 2019) used inverse graphics to capture physical quantities, while Kandukuri et al. (Kandukuri et al. 2022) enforced rigid body mechanics via convolutional encoders to identify properties like mass and friction. Yang et al. (Yang et al. 2022) and Fotiadis et al. (Fotiadis et al. 2023) utilized encoder-decoder models for low-dimensional physical coordinates. Unlike our approach, these methods depend on prior knowledge or known equations, lacking the ability to infer physical coordinates and governing equations from unfamiliar videos.
Multimodal large language models on vision tasks
@@
Recent research on multimodal large language models in vision tasks has demonstrated significant advancements. Wu et al. (Wu et al. 2025) meticulously designed a reasoning chain to unlock the potential of MLLMs in object detection tasks. Liu et al. (Liu et al. 2025) further combined physics engines with MLLMs, enabling MLLMs to analyze motion targets and their physical properties from images to generate physics-grounded videos. Additionally, extensive research has focused on fine-tuning or reasoning chains to enhance MLLMs’ video understanding capabilities (Zhou et al. 2024; Li et al. 2024; Zeng et al. 2024). In contrast to these efforts, we pioneer the exploration of unlocking MLLMs’ ability to discover equations from videos through carefully designed tools and reasoning chains.
Visual symbolic regression
@@
High-dimensional symbolic regression methods for videos is a novel scientific task that typically follows two paradigms: (1) object tracking and equation discovery, where trajectories are extracted and equations identified (Luan, Liu, and Sun 2022; Zhang, Liu, and Sun 2024); (2) encoder-decoder models for latent physical coordinates, mapping data into low-dimensional spaces for equation regression (Udrescu and Tegmark 2021; Champion et al. 2019). Our approach uses MLLMs’ visual understanding and scientific knowledge for zero-shot discovery of both coordinates and equations, unifying these paradigms in one framework.
Conclusion
@@
Limitation
The performance of our framework is inherently linked to the quality of the upstream visual processing modules. Furthermore, its robustness under challenging visual conditions, such as severe occlusions or motion blur, warrants further exploration.
@@ Appendix A Data Generation
    $\partial_{t}h+\nabla\cdot h\mathbf{u}$ =0=0,= 0 , (5)
@@
    We use open-source simulation data from the related work (Schmitt et al. 2024).
@@ Appendix B Additional Results
Real-world motion video
@@
Resolution $\times 500$ 500 $\times 500$ 500 $\times 500$ 500 $\times 500$ 500 $\times 500$ 500 $\times 500$ 500 $\times 100$ 100 $\times 64$ 64 $\times 128$ 128 $\times 128$ 128 $\times 128$ 128
@@
Adaptive regularization coefficient
@@
Case Method Equation Prediction
Terms
Found
False
Positives
$R^{2}$ $R^{2}@100$ $R^{2}@1000$
Linear VER Yes 0.38±0.220.38$\pm$ 0.220.38 ± 0.22 $\mathbf{9.59\pm 0.25}$ 9.68±0.209.68$\pm$ 0.209.68 ± 0.20 9.62±0.239.62$\pm$ 0.239.62 ± 0.23
$\eta$-VER Yes $\mathbf{0.35\pm 0.19}$ 9.54±0.209.54$\pm$ 0.209.54 ± 0.20 $\mathbf{9.70\pm 0.18}$ $\mathbf{9.64\pm 0.22}$
Cubic VER Yes 1.41±0.731.41$\pm$ 0.731.41 ± 0.73 9.53±0.329.53$\pm$ 0.329.53 ± 0.32 7.96±1.607.96$\pm$ 1.607.96 ± 1.60 6.16±2.786.16$\pm$ 2.786.16 ± 2.78
Circular VER Yes 0 9.91±0.079.91$\pm$ 0.079.91 ± 0.07 $\mathbf{9.99\pm 0.00}$ 9.67±0.459.67$\pm$ 0.459.67 ± 0.45
$\eta$-VER Yes 0 $\mathbf{9.92\pm 0.06}$ 9.97±0.049.97$\pm$ 0.049.97 ± 0.04 $\mathbf{9.70\pm 0.42}$
VDP VER Yes 0.80±0.400.80$\pm$ 0.400.80 ± 0.40 $\mathbf{9.69\pm 0.08}$ 9.35±0.199.35$\pm$ 0.199.35 ± 0.19 7.28±0.727.28$\pm$ 0.727.28 ± 0.72
$\eta$-VER Yes $\mathbf{0.51\pm 0.27}$ 9.45±0.059.45$\pm$ 0.059.45 ± 0.05 $\mathbf{9.52\pm 0.11}$ $\mathbf{8.31\pm 0.60}$
@@
Bayesian hyperparameter search
@@
Case Method Equation Prediction
Terms
Found
False
Positives
$R^{2}$ $R^{2}@100$ $R^{2}@1000$
Linear Bayes-SINDy Yes 0.49±0.180.49$\pm$ 0.180.49 ± 0.18 9.37±0.419.37$\pm$ 0.419.37 ± 0.41 9.42±0.359.42$\pm$ 0.359.42 ± 0.35 9.08±0.479.08$\pm$ 0.479.08 ± 0.47
Bayes-smooth Yes 2.51±1.022.51$\pm$ 1.022.51 ± 1.02 8.71±0.448.71$\pm$ 0.448.71 ± 0.44 8.56±0.548.56$\pm$ 0.548.56 ± 0.54 4.80±1.224.80$\pm$ 1.224.80 ± 1.22
Cubic Bayes-SINDy Yes 3.64±1.393.64$\pm$ 1.393.64 ± 1.39 9.09±0.209.09$\pm$ 0.209.09 ± 0.20 7.51±1.487.51$\pm$ 1.487.51 ± 1.48 3.63±2.113.63$\pm$ 2.113.63 ± 2.11
Bayes-smooth Yes 4.36±1.964.36$\pm$ 1.964.36 ± 1.96 7.72±1.047.72$\pm$ 1.047.72 ± 1.04 7.02±1.697.02$\pm$ 1.697.02 ± 1.69 2.94±1.862.94$\pm$ 1.862.94 ± 1.86
Glider Bayes-SINDy Yes 1.05±0.351.05$\pm$ 0.351.05 ± 0.35 $\mathbf{9.90\pm 0.23}$ 9.51±0.149.51$\pm$ 0.149.51 ± 0.14 8.33±0.668.33$\pm$ 0.668.33 ± 0.66
Bayes-smooth Yes 1.15±0.301.15$\pm$ 0.301.15 ± 0.30 9.15±0.149.15$\pm$ 0.149.15 ± 0.14 9.29±0.099.29$\pm$ 0.099.29 ± 0.09 7.89±0.267.89$\pm$ 0.267.89 ± 0.26
VER Yes $\mathbf{0.84\pm 0.17}$ 9.78±0.089.78$\pm$ 0.089.78 ± 0.08 $\mathbf{9.90\pm 0.07}$ $\mathbf{9.23\pm 0.18}$
@@
Time cost
@@ Appendix C Supplementary Materials
Bruss AE-SINDy $\mathbf{8.62\pm 0.23}$ 7.72±0.447.72$\pm$ 0.447.72 ± 0.44 6.91±0.816.91$\pm$ 0.816.91 ± 0.81 6.63±0.926.63$\pm$ 0.926.63 ± 0.92
VER 8.50±0.078.50$\pm$ 0.078.50 ± 0.07 $\mathbf{8.32\pm 0.18}$ $\mathbf{7.68\pm 0.61}$ $\mathbf{6.96\pm 0.82}$
Water AE-SINDy 9.64±0.089.64$\pm$ 0.089.64 ± 0.08 9.00±0.359.00$\pm$ 0.359.00 ± 0.35 8.07±0.558.07$\pm$ 0.558.07 ± 0.55 7.03±1.307.03$\pm$ 1.307.03 ± 1.30